OpenStudy (anonymous):

How do I set up this proportion? If it costs $12 to buy 5 pounds of grapes, how much will you have to pay to buy 9 pounds?

OpenStudy (amistre64):

12 is to 5, as "what" is to 9?

OpenStudy (amistre64):

12:5 :: x:9

OpenStudy (anonymous):

Can you explain though. I am having trouble with proportions.

OpenStudy (amistre64):

do you know how to cross multiply?

OpenStudy (anonymous):

Yes.

OpenStudy (amistre64):

thats all this is ... \[\frac{12}5=\frac x9\]

OpenStudy (anonymous):

That is how you set up the proportion?

OpenStudy (amistre64):

yep notice that 12 is to 5, as x is to 9

OpenStudy (anonymous):

So, Kelly makes $80 for working a 3-hour shift at a restaurant. If she made $210, how many hours did she work? 80/3 = 210/x ?